# Fine-tuning foundational models to code diagnoses from veterinary health records

### v1

@@ 1.1.3 Veterinary Clinical LLMs
Much prior work has focused on automated clinical coding in human EHRs [dong2022automated, feng2024automated, si2021deep, van2023diagnosis, venkataraman2020fastag, wang2023optimized, kavuluru2015empirical] for insurance claims and other higher level tasks such as enhancing patient outcomes [hamaenhancing]. However, limited clinical coding work has been done in the veterinary space with the lack of a similar financial incentive [nie2018deeptag, zhang2019vettag, jiang2023vetllm, farrell2023petbert, hur2020domain]. While not publicly available, DeepTag [nie2018deeptag] and VetTag [zhang2019vettag] both served as strong inspirations for this study. Two other foundational veterinary clinical LLMs that are publicly available and were considered and tested for this study are VetBERT (2020) [hur2020domain] and PetBERT (2023) [farrell2023petbert].
@@
VetBERT aimed to identify the reason or indication for a given drug to aid in antimicrobial stewardship. The VetCompass Australia corpus containing over 15 million clinical records from 180 veterinary clinics in Australia (5% of all vet clinics there) [mcgreevy2017vetcompass] was used for pre-training. A portion of these records were manually mapped to 52 codes in VeNom (Veterinary Nomenclature) [venomcoding] for fine-tuning. VetBERT was based on the ClinicalBERT [huang2019clinicalbert] foundational LLM and achieved strong predictive accuracy for the downstream task of identification of reason for drug administration.
@@
Most recently, Farrell et al. created PetBERT, a BERT-based LLM trained on over 5 million EHRs from the Small Animal Veterinary Surveillance Network (SAVSNET) in the United Kingdom with 500 million words. The focus was on first-opinion EHRs or primary care clinics that “often lack equivalent diagnostic certainty to referral clinics, frequently drawing upon specialised expertise and a wealth of supplementary diagnostic information” [farrell2023petbert]. Further, PetBERT-ICD was created for the specific clinical coding task of classifying EHRs to 20 ICD-11 [icd11] diagnosis categories, using 8,500 manually annotated records (7,500 training, 1,000 final evaluation). It achieved a weighted average F1 score of 84 in this task [farrell2023petbert].
1.1.4 Non-Clinical LLMs
@@
Non-clinical pre-trained LLMs are relatively more abundant than those in the clinical domain due to less privacy concerns (e.g., HIPAA). Several prominent and tractably sized non-clinical LLMs considered and tested for this study were BERT (2018) [bert2018], RoBERTa (2019) [DBLP:journals/corr/abs-1907-11692], and GPT-2 (2019) [radford2019language].
@@
Bidirectional Encoder Representations from Transformers (BERT) was introduced by Google in 2018 as a novel encoder-only variation on the original transformer model [vaswani2017transformer]. It differs from previous transformer models by using masked language modeling (MLM) and next sentence prediction (NSP) pre-training and by removing the decoder component of the transformer. This procedure makes the model specifically designed for problems involving fine-tuning on a downstream task. Several other models discussed in this paper (i.e., MedicalAI ClinicalBERT, VetBERT, and PetBERT) were based on the BERT architecture. BERT was pre-trained on a corpus of 800 million words from BooksCorpus [zhu2015] and 2.5 billion words from Wikipedia. It was benchmarked on eleven standard NLP tasks and achieved state-of-the-art results for its time.
@@
It was later determined that the foundational BERT LLM was “significantly undertrained;” RoBERTa [DBLP:journals/corr/abs-1907-11692] was offered as a better-constructed alternative. An enhanced pre-training dataset was assembled using the BooksCorpus and Wikipedia datasets used to train BERT as well as CC-News - 63 million scraped news articles from 2016 to 2019, OpenWebText [peterson2019] - scraped text from web links shared on Reddit, and Stories [trinh2018simple] - a subset of the CommonCrawl [nagel2016common] foundational dataset. With this enhanced dataset and an improved training scheme, RoBERTa achieved state-of-the-art performance on a variety of standard NLP benchmarks [DBLP:journals/corr/abs-1907-11692].
@@
In contrast to BERT, Generative Pre-trained Transformer 2 (GPT-2) is a decoder-only transformer released by OpenAI. The decoder-only architecture relies on unidirectional pre-training and was designed for a variety of zero-shot tasks. GPT-2 was pre-trained on an OpenAI-generated WebText dataset consisting of over eight million documents scraped from the web in December 2017. Text from Wikipedia was specifically excluded from WebText due to its common usage in other foundational models. It achieved state-of-the-art results on seven standard NLP benchmarks in a zero-shot setting.
@@ 1.2 Aims
This study aimed to extend and improve upon prior work in the field of automated clinical coding for veterinary EHRs to facilitate standardization of records, veterinary patient health and research, and the creation of data linkages to support One Health approaches to problem solving. Available manually coded data from the CSU VTH was used to determine if fine-tuning existing foundational models could achieve state-of-the-art results for automated veterinary clinical coding to 7,739 SNOMED-CT diagnosis codes, the largest set of diagnosis codes yet used in a veterinary context. It was found that fine-tuning the foundational model GatorTron performed the best with an average weighted F1 score of 74.9 and an exact match rate of 51.6%.
@@ 2 Materials and Methods
Figure 1: Colorado State University (CSU) automated clinical coding pipeline from raw records to a standardized Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM) instance, with future work noted. VTH = Veterinary Teaching Hospital; LLM = large language model.
@@
Figure 2: Sample excerpt from veterinary medical summary documents from the Colorado State University Veterinary Teaching Hospital electronic health records. The majority of fine-tuned models, with varying degrees of confidence, were able to correctly identify the appropriate code as the most likely applicable code to this record. This is in spite of the fact that none of the key words from the code description were found in the record text.
@@ 2.1 Data
SNOMED-CT terminologies are version controlled with older codes being occasionally upgraded or deprecated in favor of new “active” codes. The version of SNOMED-CT used here was downloaded from OHDSI in December 2022. For this study, both active and inactive codes were allowed to capture all available data. The set of diagnosis codes used for manual annotation over many years at the CSU VTH naturally diverged somewhat from being included and active in the latest SNOMED-CT release and includes some (530, 7%) currently inactive codes. There are inactive-to-active mappings available for many codes that can translate outdated vocabulary sets to conform with current standards when needed [historicalassociation].
@@ 2.1.2 Input Fields and Preprocessing
To understand the contributions of two different sections of the medical summary on model performance, a basic analysis was conducted. In consultation with the CSU VTH Medical Records team, it was determined that the diagnosis section contained the primary information used to inform diagnosis coding. Additionally, the longer narrative assessment section was included for the broad context it provided to visits. Other sections of the EHR were not explored in detail due to model restrictions on input text length.
The final input text for this task was a concatenation of diagnosis and assessment sections, in that order. Minimal text cleaning was performed (only simple replacement of Hypertext Markup Language (HTML) characters, lowercase, etc.), and no de-identification methods were used. Records were tokenized according to the custom tokenizers associated with each pre-trained LLM and truncated to the corresponding maximum length (truncation rate did not exceed 23% for any tokenizer, and records contained 284 tokens on average). Target codes for each record were then one-hot encoded for training. Using a stratified sampling technique to account for class imbalances, the full fine-tuning dataset was split into training (80%), validation (10%), and test (10%) sets.
@@ 2.2 Evaluation Metrics
These metrics were chosen in accordance with those reported by the VetTag team and as standard multi-label classification metrics where a skewed class distribution is present [resnik2010evaluation].
@@ 2.3 Models
Table 1: Selected clinical language models characterized by pre-training data source and volume, size (number of parameters), number of transformer blocks, number of self-attention heads, and embedding dimension. All models were fine-tuned on Colorado State University Veterinary Teaching Hospital records. M = million; B = billion; UK = United Kingdom; Vet = veterinary; EHRs = electronic health records; BERT = bidirectional encoder representations from transformers; GPT = generative pre-trained transformer; PSVG = private specialty veterinary group.
*Model not fine-tuned for this project and utilized a hierarchical training objective: shown only for comparison.
Existing state-of-the-art human clinical LLMs (GatorTron, MedicalAI ClinicalBERT, medAlpaca) and veterinary (VetBERT, PetBERT) and general knowledge LLMs (BERT, BERT Large, RoBERTa, GPT-2, and GPT-2 XL) were leveraged by fine-tuning them for veterinary clinical coding to 7,739 SNOMED-CT diagnoses (Table 1). All models were fine-tuned under a consistent training framework that included a batch size of 32, a binary cross-entropy loss function, an AdamW optimizer, scheduled optimization with 5,000 warmup steps and a plateau learning rate (constant learning rate after warmup) of $3*10^{-5}$, and an early stopping criterion with five epoch patience (number of epochs without improved results required to trigger an early stop). All models’ pooler layers were allowed to update, and dropout (rate of 0.25) and classifier layers were appended to each. Fine-tuning was performed on as many transformer blocks as possible, given computational constraints. The results for all fine-tuned models were compared to the results of the current state-of-the-art model in veterinary clinical coding, VetTag. Additionally, all models were tested “out-of-the-box” with all pre-trained weights frozen except for pooler and classifier layers, which were allowed to be fine-tuned. Unlike DeepTag and VetTag, a hierarchical training objective [zhang2019vettag] was not used in this study.
@@ 3 Results
State-of-the-art results were achieved on veterinary clinical coding to the largest set of diagnosis concepts (Table 2). GatorTron fine-tuned for this task achieved the best results with an average weighted F1 score of 74.9 and an exact match rate of 51.6% on a held-out test set (10% of the labeled CSU VTH data). Many smaller models considered for this study (e.g., GPT-2 and MedicalAI ClinicalBERT) still performed comparably to the larger GatorTron (Table 2). Overall, all models had higher precision compared to recall.
All models were also tested “out-of-the-box” with all pre-trained weights frozen during fine-tuning (Table 3). On average, models’ F1 scores were 48.4 points lower in the “out-of-the-box” setting compared with more expansive fine-tuning. Particularly of note were the BERT models which failed to converge to a reasonable solution and the GPT-2 models which performed strongly relative to all other models.
@@
GatorTron 74.9$\pm$0.29 80.8$\pm$1.14 71.8$\pm$1.00 51.6$\pm$0.29 21.9$\pm$2.87 hrs
MedicalAI ClinicalBERT 68.7$\pm$0.57 78.0$\pm$1.15 63.9$\pm$0.29 45.2$\pm$0.86 1.5$\pm$0.14 hrs
medAlpaca 67.1$\pm$0.90 79.1$\pm$0.66 61.7$\pm$1.00 41.6$\pm$1.46 14.1$\pm$0.14 hrs
VetBERT 69.5$\pm$0.87 78.7$\pm$0.86 64.7$\pm$0.90 46.5$\pm$0.66 2.9$\pm$0.38 hrs
PetBERT 69.4$\pm$2.00 77.6$\pm$1.88 65.3$\pm$2.17 46.4$\pm$1.41 2.7$\pm$0.50 hrs
BERT Base 68.5$\pm$1.93 77.5$\pm$0.57 63.9$\pm$3.19 45.8$\pm$1.27 3.6$\pm$0.57 hrs
BERT Large 70.4$\pm$0.14 78.4$\pm$0.29 66.2$\pm$0.38 47.2$\pm$0.57 7.4$\pm$1.15 hrs
RoBERTa 67.7$\pm$4.88 76.5$\pm$3.02 63.3$\pm$5.41 44.6$\pm$4.81 2.6$\pm$0.87 hrs
GPT-2 Small 68.3$\pm$1.51 78.3$\pm$1.17 63.4$\pm$1.74 44.6$\pm$1.08 4.1$\pm$0.87 hrs
GPT-2 XL 71.7$\pm$1.62 80.4$\pm$0.76 67.2$\pm$2.01 47.8$\pm$1.74 15.2$\pm$1.58 hrs
@@
Table 2: Results computed on electronic health records text test sets. Only diagnosis and assessment fields from the medical summary were used. All Systematized Nomencalture of Medicine - Clinical Terms (SNOMED-CT) diagnosis codes that were manually assigned for these visits were used regardless of frequency or standard status. Training was not supplemented using the SNOMED-CT hierarchy. Testing was performed three times for each model to allow for the computation of 95% confidence intervals ($\mu\pm t_{.975,2}*\sigma/\sqrt{3}$). Test sets were standardized across models through set random seeds.
@@
Table 3: Out-of-the-box results computed on electronic health records text test sets. Foundational model weights were frozen and only pooler and classifier layers were allowed to update during the fine-tuning process. All other training practices were the same as those in Table 1.

### v2

@@ 1.1.3 Veterinary Clinical LMs
Much prior work has focused on automated clinical coding in human EHRs [69, 72, 105, 112, 114, 116, 84] for insurance claims and other higher level tasks such as enhancing patient outcomes [74]. However, limited clinical coding work has been done in the veterinary space with the lack of a similar financial incentive [94, 121, 81, 71, 79]. While not publicly available, DeepTag [94] and VetTag [121] both served as strong inspirations for this study. Two other foundational veterinary clinical LMs that are publicly available and were considered and tested for this study are VetBERT (2020) [79] and PetBERT (2023) [71].
@@
VetBERT aimed to identify the reason or indication for a given drug to aid in antimicrobial stewardship. The VetCompass Australia corpus containing over 15 million clinical records from 180 veterinary clinics in Australia (5% of all vet clinics there) [91] was used for pre-training. A portion of these records were manually mapped to 52 codes in VeNom (Veterinary Nomenclature) [115] for fine-tuning. VetBERT was based on the ClinicalBERT [78] foundational LM and achieved strong predictive accuracy for the downstream task of identification of reason for drug administration.
@@
Most recently, Farrell et al. created PetBERT, a BERT-based LM trained on over five million EHRs from the Small Animal Veterinary Surveillance Network (SAVSNET) in the United Kingdom with 500 million words. The focus was on first-opinion EHRs or primary care clinics that “often lack equivalent diagnostic certainty to referral clinics, frequently drawing upon specialised expertise and a wealth of supplementary diagnostic information” [71]. Further, PetBERT-ICD was created for the specific clinical coding task of classifying EHRs to 20 ICD-11 [80] diagnosis categories, using 8,500 manually annotated records (7,500 training, 1,000 final evaluation). It achieved a weighted average F1 score of 84 in this task [71].
1.1.4 Non-Clinical LMs
@@
Non-clinical pre-trained LMs are relatively more abundant than those in the clinical domain due to less privacy concerns (e.g., HIPAA). Several prominent and tractably sized non-clinical LMs considered and tested for this study were BERT (2018) [68], RoBERTa (2019) [87], GPT-2 (2019) [103], DeBERTa V3 (2021) [76], and ModernBERT (2024) [117].
@@
Bidirectional Encoder Representations from Transformers (BERT) was introduced by Google in 2018 as a novel encoder-only variation on the original transformer model [113]. It differs from previous transformer models by using masked language modeling (MLM) and next sentence prediction (NSP) pre-training, and by removing the decoder component of the transformer. This procedure makes the model specifically designed for problems involving fine-tuning on a downstream task. Several other models discussed in this paper (MedicalAI ClinicalBERT, VetBERT, and PetBERT) were based on the BERT architecture. BERT was pre-trained on a corpus of 800 million words from BooksCorpus [122] and 2.5 billion words from Wikipedia. It was benchmarked on eleven standard NLP tasks and achieved state-of-the-art results for its time.
@@
It was later determined that the foundational BERT LM was “significantly undertrained;” RoBERTa [87] was offered as a better-constructed alternative. An enhanced pre-training dataset was assembled using the BooksCorpus and Wikipedia datasets used to train BERT as well as CC-News - 63 million scraped news articles from 2016 to 2019, OpenWebText [101] - scraped text from URLs shared on Reddit, and Stories [111] - a subset of the CommonCrawl [93] foundational dataset. With this enhanced dataset and an improved training scheme, RoBERTa achieved state-of-the-art performance on a variety of standard NLP benchmarks [87].
The foundational BERT and RoBERTa LMs were further improved upon by the introduction of Decoding-enhanced BERT with disentangled attention (DeBERTa) [76] and ModernBERT [117]. DeBERTa utilized a disentangled attention mechanism, an enhanced mask decoder, and a virtual adversarial training method to improve model efficiency and performance. It was trained on an English Wikipedia dump, BookCorpus [122], OpenWebText [101], and Stories [111]. ModernBERT utilized rotary positional embeddings (RoPE), unpadding, and flash attention to improve model efficiency. Both model architectures demonstrated superior performance relative to previous BERT-based models.
@@
In contrast to BERT, Generative Pre-trained Transformer 2 (GPT-2) [103] is a decoder-only transformer released by OpenAI. The decoder-only architecture relies on unidirectional pre-training and was designed for a variety of zero-shot tasks. GPT-2 was pre-trained on an OpenAI-generated WebText dataset consisting of over eight million documents scraped from the web in December 2017. Text from Wikipedia was specifically excluded from WebText due to its common usage in other foundational models. It achieved state-of-the-art results on seven standard NLP benchmarks in a zero-shot setting.
@@ 1.2 Aims
This study aimed to extend and improve upon prior work in the field of automated clinical coding for veterinary EHRs to facilitate standardization of records, veterinary patient health and research, and the creation of data linkages to support One Health approaches to problem solving. Available manually coded data from the CSU VTH was used to determine if fine-tuning existing foundational models could achieve state-of-the-art results for automated veterinary clinical coding to 7,739 SNOMED-CT diagnosis codes, the largest set of diagnosis codes yet used in a veterinary context. It was found that fine-tuning the foundational model GatorTron performed the best with an average weighted F1 score of 76.9 and an exact match rate of 52.2%.
@@ 2 Materials and Methods
Figure 1: Colorado State University (CSU) automated clinical coding pipeline from raw records to a standardized Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM) instance, with future work noted. VTH = Veterinary Teaching Hospital; LM = language model.
@@
Figure 2: Sample excerpt from veterinary medical summary documents from the Colorado State University Veterinary Teaching Hospital electronic health records.
@@ 2.1 Data
SNOMED-CT terminologies are version controlled with older codes being occasionally upgraded or deprecated in favor of new “active” codes. The version of SNOMED-CT used here was downloaded from OHDSI in December 2022. For this study, both active and inactive codes were allowed for practical reasons. The set of diagnosis codes used for manual annotation over many years at the CSU VTH naturally diverged somewhat from being included and active in the latest SNOMED-CT release and includes some (530, 7%) currently inactive codes. There are inactive-to-active mappings available for many codes that can translate outdated vocabulary sets to conform with current standards when needed [77].
@@ 2.1.2 Input Fields and Preprocessing
In consultation with the CSU VTH Medical Records team, it was determined that the diagnosis section contained the primary information used to inform diagnosis coding. This finding was confirmed by a sensitivity analysis early on (results not shown) that examined model performance for each section and with different combinations of medical summary sections (i.e., diagnosis, assessment, presenting complaint, history, physical exam, procedures, and treatments) used as input text. The only section that independently drove model performance was diagnosis. Other sections did not exhibit the capacity for strong independent performance, and their inclusion in the input text alongside diagnosis did not significantly improve performance. Only the longer narrative assessment section was appended to diagnosis for the broad context it provided to visit summaries. Other sections of the EHR were not included due to model restrictions on input text length and in the interest of minimizing text truncation.
Based on the CSU VTH Medical Records team and the early on sensitivity analysis, the final input text for this task was a concatenation of diagnosis (including histopathology diagnosis text) and assessment sections, in that order. Minimal text cleaning was performed (only simple replacement of Hypertext Markup Language (HTML) characters, lowercase, etc.), and no de-identification methods were used. Records were tokenized according to the custom tokenizers associated with each pre-trained LM and truncated to the corresponding maximum length (truncation rate did not exceed 23% for any tokenizer, records contained 260 tokens on average, and the rate of unknown tokens used did not exceed 0.36% for any tokenizer). Target codes for each record were then one-hot encoded for training. Using a stratified sampling technique to account for class imbalances, the full fine-tuning dataset was split into training (80%), validation (10%), and test (10%) sets.
@@ 2.2 Evaluation Metrics
These metrics were chosen in accordance with those reported by the VetTag team and as standard multi-label classification metrics where a skewed class distribution is present [104].
@@ 2.3 Models
Table 1: Selected clinical language models characterized by pre-training data source and volume, size (number of parameters), number of transformer blocks, number of self-attention heads, and embedding dimension. All models were fine-tuned on Colorado State University Veterinary Teaching Hospital records. M = million; B = billion; UK = United Kingdom; Vet = veterinary; EHRs = electronic health records; BERT = bidirectional encoder representations from transformers; GPT = generative pre-trained transformer; PSVG = private specialty veterinary group. *Model not fine-tuned for this project and utilized a hierarchical training objective: shown only for comparison.
@@
DeBERTa V3 Wikipedia, BookCorpus, OpenWebText, Stories 184M 12 12 768
ModernBERT Web documents, code, scientific literature (2T tokens) 149M 22 12 768
Clinical ModernBERT MIMIC-IV, PubMed, Clinical Codes (e.g., ICD) (13B tokens) 136M 22 12 768
@@
Existing state-of-the-art human clinical LMs (GatorTron, MedicalAI ClinicalBERT, medAlpaca, Clinical ModernBERT), veterinary (VetBERT, PetBERT), and general knowledge LMs (BERT, BERT Large, RoBERTa, GPT-2, GPT-2 XL, DeBERTa V3, ModernBERT) were leveraged by fine-tuning them for veterinary clinical coding to 7,739 SNOMED-CT diagnoses (Table LABEL:table:Models). Ensemble, mixture of experts, and agentic models were not considered for this study. All models were fine-tuned under a consistent training framework (except for Bert Large, which did not feature a dropout layer) that included a batch size of 32, a binary cross-entropy loss function, an AdamW optimizer, scheduled optimization with 5,000 warmup steps and a plateau learning rate (constant learning rate after warmup) of $3*10^{-5}$, and an early stopping criterion with five epoch patience and a maximum of 50 epochs (number of epochs without improved results required to trigger an early stop). Through experimentation, it was determined that best model performance was achieved when the maximum number of transformer blocks were allowed to update in the fine-tuning process. All transformer blocks for all models were fine-tuned with the exception of GatorTron and medAlpaca, which were too large to fully fine-tune given the computational resource available. No embeddings were fine-tuned. Custom pooler layers consisting of a fully-connected linear layer followed by a Tanh activation function, dropout layers (rate of 0.25), and linear classifier layers were appended to each model. No gradient accumulation or clipping, or progressive unfreezing methods were used. No additional pretraining was performed to isolate the effects of fine-tuning on model performance. The results for all fine-tuned models were compared to the results of the current state-of-the-art model in veterinary clinical coding, VetTag. Additionally, all models were tested “out-of-the-box” to determine if it was necessary to fine-tune the models for this task. In this case, all pre-trained weights were frozen and pooler and classifier layers were fine-tuned under the same conditions described above. Unlike DeepTag and VetTag, a hierarchical training objective [121] was not used in this study.
@@ 3 Results
State-of-the-art results were achieved on veterinary clinical coding to the largest set of diagnosis concepts (Table 2). GatorTron fine-tuned for this task achieved the best results with an average weighted F1 score of 76.9 and an exact match rate of 52.2% on a held-out test set (10% of the labeled CSU VTH data). Many smaller models considered for this study (e.g., VetBERT and RoBERTa) still performed comparably to the larger GatorTron (Table 2). Overall, all models had higher precision compared to recall.
The “out-of-the-box” results were poor (Table 3), confirming the need for fine-tuning. On average, models’ F1 scores were 52.9 points lower in the “out-of-the-box” setting compared with more expansive fine-tuning. Particularly of note were the BERT and DeBERTa models which failed to converge to a reasonable solution and the GPT-2 models which performed strongly relative to all other models. These results are expected as adapter-based fine-tuning can result in unstable model performance due to a failure to “capture complex data patterns” [102].
@@
[…]
@@
Table 2: Results computed on electronic health record text test sets. Only diagnosis and assessment fields from the medical summary were used based on the CSU VTH Medical Record coders and an early on sensitivity analysis. All manually assigned Systematized Nomenclature of Medicine - Clinical Terms (SNOMED-CT) diagnosis codes for these visits were used regardless of frequency or standard status. Training was not supplemented using the SNOMED-CT hierarchy. Testing was performed three times for each model to allow for the computation of 95% confidence intervals ($\mu\pm t_{.975,2}*\sigma/\sqrt{3}$). Test sets were standardized across models through set random seeds.
@@
DeBERTa V3 1.1 6.3 0.6 5.2
ModernBERT 7.4 23.4 5.9 11.9
Clinical ModernBERT 23.1 51.0 17.2 16.0
Table 3: Out-of-the-box results computed on electronic health record text test sets. Foundational model weights were frozen and only pooler and classifier layers were allowed to update during the fine-tuning process. All other training practices were the same as those in Table 1.